# Pairing mechanism and superconductivity in pressurized La5Ni3O11

### v1

@@ Results
Here, $U$, $V$, and $J_{H}$ denote the intra-orbital, inter-orbital Hubbard repulsion, and the coupling of Hund (and the pair hopping) respectively, which satisfy the relation $U=V+2J_{H}$. $\tilde{\mu}\in(1,2)$ labels the two orbitals ($d_{z^{2}}$ and $d_{x^{2}-y^{2}}$) associated with each Ni atom. $i$ denotes the lattice sites belonging to the bilayer subsystem. We employ the multi-orbital RPA approach  takimoto2004strong ; yada2005origin ; kubo2007pairing ; graser2009near ; liu2013d+ ; zhang2022lifshitz ; kuroki101unconventional to treat this Hamiltonian. By renormalization, the spin susceptibility $\bm{\chi}^{(s)}$ and charge susceptibility $\bm{\chi}^{(c)}$ can be defined as Eq. (Pairing mechanism and superconductivity in pressurized La5Ni3O11):
@@
Where $\bm{\chi}^{(0)}$ is bare susceptibility for the non-interacting case, expressed as a tensor $\bm{\chi}^{(0)pq}_{st}$, with $p,q,s,t$ as orbital indices. Similarly, $\bm{\chi}^{(s)}$ can be expressed as $\bm{\chi}^{(s)pq}_{st}$. $U^{(s,c)}$ is the renormalized interaction strength, which is represented as a $4^{2}\times 4^{2}$ matrix in the bilayer subsystem. Note that there is a critical interaction strength $U_{c}^{(s/c)}$ for the spin/charge susceptibility. When $U\geq U_{c}^{(s,c)}$, the denominator matrix in Eq. (Pairing mechanism and superconductivity in pressurized La5Ni3O11) will have zero eigenvalues for certain values of $\bm{k}$, causing the renormalized spin or charge susceptibility to diverge. This divergence indicates the onset of magnetic or charge order. Generally, repulsive Hubbard interactions suppress the charge susceptibility, but enhance the spin susceptibility  takimoto2004strong ; yada2005origin ; kubo2007pairing ; graser2009near ; liu2013d+ ; zhang2022lifshitz ; kuroki101unconventional . Therefore, $U_{c}^{(s)}<U_{c}^{(c)}$ and we denote $U_{c}^{(s)}$ as $U_{c}$ in the bilayer subsystem in the following. When fixing $J_{H}=U/6$, we find $U_{c}\approx 1.15$ eV.
Defining $\chi^{s}(\bm{q})$ as the maximum eigenvalue of $\bm{\chi}^{(s)}$ at each momentum, Fig. 4(b) shows its distribution in the brillouin zone for $U=1$ eV$<U_{c}$ of the bilayer subsystem. Notably, the distribution exhibits peaks at two unequivalent momenta, which we have labeled as $\mathbf{Q}_{1}$ and $\mathbf{Q}_{2}$. These two momenta precisely correspond to the two FS nesting vectors of the bilayer subsystem, as illustrated in Fig. 4(a), and the vector $\mathbf{Q}_{1}=(\pi,0)$ is associated with the highest intensity of spin susceptibility, which is very similar to the case of the La3Ni2O7 superconductor under pressure WangQH2023 ; YangF2023 .
When $U<U_{c}$, the spin fluctuations can mediate SC in the bilayer subsystem, whose $T_{c}$ is related to the largest pairing eigenvalue $\lambda$ via $T_{c}\propto\omega_{D}e^{-1/\lambda}$ RPA5 , and the pairing symmetry is determined by the corresponding eigenvector (see $\bf Methods$). Fig. 4(c) illustrates the dependence of the largest pairing eigenvalue, denoted as $\lambda$, on the interaction strength $U$ for different potential pairing symmetries. The $D_{4h}$ point group of this bilayer structure allows for several possible pairing symmetries, in Fig. 4(c), we show the three leading pairing symmetries: $s$-wave, $d$-wave, and degenerate $(p_{x},p_{y})$-wave pairings. It is clear that the $s$-wave is the leading pairing symmetry and dominates other ones. The gap function of the obtained $s$-wave pairing is shown on the FS in Fig. 4(d), which displays the $s^{\pm}$ pattern. Consequently, the $\alpha$- and $\gamma$- pockets connected by the nesting vector $\mathbf{Q}_{1}$ are distributed with the strongest pairing amplitude, with their gap signs opposite. This pairing pattern is also similar to that in pressurized La3Ni2O7 luo2024high ; WangQH2023 ; YangF2023 .
@@
Table 2: Hopping parameters and site energies of TB model for La5Ni3O11 under various pressures. $x$, $z$ and $xz$ denotes hopping within and between $d_{x^{2}-y^{2}}$, $d_{z^{2}}$ orbitals, respectively. The abbreviations BI, BO, and SI represent intrabilayer, interbilayer, and single-layer in-plane hoppings, respectively. Only hoppings greater than 0.1 eV are listed here for clarity. All parameters are in unit of eV.
@@
The decreasing $\lambda\sim P$ relation caused by the reduced DOS shown in Fig. 5 parallels that in the pure La3Ni2O7 PhysRevLett.134.076001 ; huomodulation ; zhang2024structural , which closely matches the experimentally observed $T_{c}\sim P$ relation in that material li_identification_2025 . However, it seems at a glance that such a decreasing $\lambda\sim P$ relation conflicts with the $T_{c}\sim P$ relation in La5Ni3O11, which shows a pronounced increase up to around 20 GPa, followed by a gradual decline afterward shi2025superconductivityhybridruddlesdenpopperla5ni3o11 . The discrepancy between the $\lambda\sim P$ and the $T_{c}\sim P$ relations in La5Ni3O11 originates from that its alternating SL-BL structure brings into extremely weak IJC that is crucial for the establishment of the bulk SC, which can be prominently enhanced by the enhancement of $P$. In the following, we clarify this viewpoint.
Figure 6: (color online) Influence of the IJC on the pressure-dependence of the $T_{c}$. (a) Schematic of the intrinsic S–N–S Josephson Junction structure of La5Ni3O11, where the superconducting (S) layers correspond to the La3Ni2O7 bilayer and the normal (N) layers correspond to the La2NiO4 single-layer. (b) Schematic of the interlayer distance $d_{\perp}$ dependence of the hopping integral $t_{z}^{BS}$ between the bilayer subsystem and the single-layer subsystem. (c) $T_{c}/\rho_{0}$ as a function of the anisotropy parameter $\eta$ that reflects the strength of the IJC. The red dashed line marks the $\eta\approx\mathrm{1}0^{-8}$ associated with the La5Ni3O11, while the blue dashed line marks the $\eta\approx\mathrm{1}0^{-4}$ associated with the La3Ni2O7. (d) Pressure-dependence of the superconducting $T_{c}$ (in unit of $\omega_{D}$) in La5Ni3O11.
As shown in Fig. 6(a), the La5Ni3O11 hosts alternating bilayer and single-layer NiO2 subsystems, where the former supports SC and the latter is a “bad metal”. This results in an intrinsic superconductor–normal metal–superconductor (S–N–S) Josephson Junction structure. In this structure, the establishment of bulk SC requires not only intra-bilayer pairing but also inter-bilayer phase coherence. Physically, the inter-bilayer phase coherence is established via combined tunneling of a Cooper pair from a superconducting bilayer to its adjacent one, i.e. the IJC, across the intervening normal-metallic single-layer. This is a second-order process of the inter-bilayer single-particle tunneling, and hence a fourth-order process of the weak bilayer-single-layer tunneling for a single particle. As such an IJC is extremely weak, it will experience a prominent enhancement with the increase of the pressure. In the following, we provide a more quantitative clarification for this point.
@@
In a quasi-2D layered system such as La5Ni3O11 and La3Ni2O7, the bulk $T_{c}$ is related to the anisotropy parameter $\eta$ through the following relation in the limit of $\eta\to 0$ kopec2000superconducting ,
@@
Here $\rho_{0}$ indicates the intralayer phase stiffness reflecting the pairing within the bilayer subsystem, which is approximated as the pairing temperature obtained from the RPA calculation, i.e. $\omega_{D}e^{-1/\lambda}$. Here the “Debye frequency” $\omega_{D}$ represents for the typical energy scale of spin fluctuations. The anisotropy parameter $\eta$ is the ratio between the IJC strength $J_{\perp}^{BB}$ and $\rho_{0}$. Since $J_{\perp}^{BB}\propto(t_{z}^{BS})^{4}$, and $t_{z}^{BS}\sim 10^{-2}$, we have $\eta\sim 10^{-8}$ for La5Ni3O11, which is extremely weak. In contrast, for La3Ni2O7, since there exists no intervening normal-metallic single-layer to go across, we have $\eta\sim 10^{-4}$. As shown in Fig. 6(c), for La5Ni3O11, the anisotropy parameter $\eta$ locates within the regime that a slight enhancement of $\eta$ will prominently enhance the $T_{c}$; while for La3Ni2O7, $\eta$ has only a minor effect on $T_{c}$.
Under applied pressure, the strain is proportional to the pressure, implying that $\Delta d_{\perp}\propto-P$, i.e., $d_{\perp}(P)=d_{\perp}^{0}-\kappa P$, where $d_{\perp}^{0}$ is the interlayer distance at zero pressure and $\kappa$ is a proportionality constant. Therefore,
$\eta=e^{-4(d_{\perp}^{0}-\kappa P)/\xi}\equiv\eta_{0}e^{\alpha P}.$ (5)
@@
For a reasonable set of parameters, e.g. $\eta_{0}\approx 10^{-8}$ and $\alpha\approx 0.3$ GPa-1, the resulting pressure dependence of $T_{c}$ shown in Fig. 6(d) indeed exhibits a dome-shaped behavior similar to that observed in experiment shi2025superconductivityhybridruddlesdenpopperla5ni3o11 .
@@ Discussion
In this work, we investigate the electronic structure and superconducting mechanism of the hybrid RP nickelate La5Ni3O11, composed of alternating single-layer and bilayer subsystems. Based on DFT and RPA calculations, we find that SC predominantly emerges in the bilayer subsystem with an $s^{\pm}$-wave pairing symmetry, closely resembling that in pressurized La3Ni2O7. We further show that the pressure dependence of the superconducting $T_{c}$ in La5Ni3O11 is governed not only by the intralayer pairing but also by the IJC, which leads to a dome-like $T_{c}(P)$ behavior consistent with experimental observations.
@@ Methods
DFT calculations were performed by Vienna ab initio simulation package (VASP) (VASP1, ; VASP2, ), in which the projector augmented wave (PAW) (PAW1, ; PAW2, ) method with a 600 eV plane-wave cutoff is applied. The generalized gradient approximation (GGA) of Perdew-Burke-Ernzerhof (PBE) form exchange correlation potential is adopted  (PhysRevLett.77.3865, ). The convergence criterion of force was set to be 0.001 eV/Å and total energy convergence criterion was set to be $10^{-7}$ eV. A $\Gamma$-centered $19\times 19\times 5$ Monkhorst Pack k-mesh grid is used for primitive cell of $P4/mmm$ phase and a $\Gamma$-centered $14\times 14\times 5$ Monkhorst Pack k-mesh grid for primitive cell of $Cmmm$ phase. In structural relaxations, the lattice constants of $P4/mmmm$ phase we use are $a=3.77$  and $c=16.20$. DFT calculations reveal that the corresponding pressure is 12 GPa, where La5Ni3O11 enters superconducting state  shi2025superconductivityhybridruddlesdenpopperla5ni3o11 . An effective Hubbard U parameter $U=3.5$ eV was employed to account for the correlation effects of 3d electrons in Ni atoms (yang_orbital-dependent_2024, ; PhysRevB.57.1505, ). To obtain the projected TB models, we further performed Wannier downfolding as implemented in WANNIER90 (w90, ) package, in which the good convergences were reached.
@@
×$\left.\times c_{s}^{\dagger}(\bm{k}_{2}+\bm{k},0)c_{t}(\bm{k}_{2},0)\right\rangle_{0},$ (7)
@@
$U^{(s)pq}_{st}=\left\{\begin{aligned} U,~{}~{}~{}p=q=s=t\\ J_{H},~{}~{}~{}p=q\neq s=t\\ J_{H},~{}~{}~{}p=s\neq q=t\\ V,~{}~{}~{}p=t\neq s=q\\ \end{aligned}\right.$
$U^{(c)pq}_{st}=\left\{\begin{aligned} U,~{}~{}~{}p=q=s=t\\ 2V-J_{H},~{}~{}~{}p=q\neq s=t\\ J_{H},~{}~{}~{}p=s\neq q=t\\ 2J_{H}-V,~{}~{}~{}p=t\neq s=q\\ \end{aligned}\right.$
@@
Note that there is a critical interaction strength $U_{c}^{(s,c)}$ for spin and charge respectively. When $U>U_{c}^{(s,c)}$, the denominator matrix in Eq. Pairing mechanism and superconductivity in pressurized La5Ni3O11 will have zero eigenvalues for some $\mathrm{q}$ and the renormalized spin or charge susceptibility diverges there, which invalidates the RPA treatment RPA1 ; RPA2 ; RPA3 ; RPA4 ; RPA5 ; RPA6 ; RPA7 ; RPA8 ; PhysRevB.110.L180501 . This divergence of spin susceptibility for $U>U_{c}^{(s)}$ implies magnetic order, while that of the charge susceptibility for $U>U_{c}^{(c)}$ implies charge order.
@@
Here $\alpha$/$\beta$ is band indice. And effective pairing interaction vertex $V^{\alpha\beta}(\mathbf{k},\mathbf{q})$ has the form:
@@
Here $\xi^{\alpha}\left(\mathbf{k}\right)$ are the eigenvector corresponding to $\alpha$-th eigenvalue (relative to the chemical potential $\mu_{c}$) of the TB Hamiltonian matrices. Within the mean-field approximation, one can derive a self-consistent equation for the pairing gap, which when linearized near $T_{c}$ RPA1 ; RPA2 ; RPA3 ; RPA4 ; RPA5 ; RPA6 ; RPA7 ; RPA8 ; PhysRevB.110.L180501 yields the linearized gap equation (12):
@@
Here $\beta_{c}$=$\frac{1}{\mathrm{k_{B}}T_{c}}$ represents the critical temperature of SC. Choosing a thin energy shell near the Fermi level, Eq. (12) becomes the eigenvalue problem of the effective interaction matrix $V^{\alpha\beta}$, which determines the $T_{c}$ and the leading pairing symmetry of the system. To be specific, $\Delta_{\alpha}(\mathbf{k})$ represents the relative gap function on the $\alpha$-th FS patches near $T_{c}$, and eigenvalue $\lambda$ is related to $T_{c}$ through $\lambda^{-1}=\mathrm{ln}(1.13\frac{\hbar\omega_{D}}{k_{\mathrm{B}}}T_{c})$. The leading pairing symmetry is determined by the largest eigenvalue $\lambda$ of Eq. (13):
[…]

### v2

@@ Results
Here, $U$, $V$, and $J_{H}$ denote the intra-orbital, inter-orbital Hubbard repulsion, and the coupling of Hund (and the pair hopping) respectively, which satisfy the relation $U=V+2J_{H}$. $\tilde{\mu}\in(1,2)$ labels the two orbitals ($d_{z^{2}}$ and $d_{x^{2}-y^{2}}$) associated with each Ni atom. $i$ denotes the lattice sites belonging to the bilayer subsystem. We employ the multi-orbital RPA approach  Takimoto et al. (2004); Yada and Kontani (2005); Kubo (2007); Graser et al. (2009); Liu et al. (2013); Zhang et al. (2022); Kuroki et al. (2008) to treat this Hamiltonian. By renormalization, the spin susceptibility $\bm{\chi}^{(s)}$ and charge susceptibility $\bm{\chi}^{(c)}$ can be defined as Eq. (Pairing mechanism and superconductivity in pressurized La5Ni3O11):
@@
Where $\bm{\chi}^{(0)}$ is bare susceptibility for the non-interacting case, expressed as a tensor $\bm{\chi}^{(0)pq}_{st}$, with $p,q,s,t$ as orbital indices. Similarly, $\bm{\chi}^{(s)}$ can be expressed as $\bm{\chi}^{(s)pq}_{st}$. $U^{(s,c)}$ is the renormalized interaction strength, which is represented as a $4^{2}\times 4^{2}$ matrix in the bilayer subsystem. Note that there is a critical interaction strength $U_{c}^{(s/c)}$ for the spin/charge susceptibility. When $U\geq U_{c}^{(s,c)}$, the denominator matrix in Eq. (Pairing mechanism and superconductivity in pressurized La5Ni3O11) will have zero eigenvalues for certain values of $\bm{k}$, causing the renormalized spin or charge susceptibility to diverge. This divergence indicates the onset of magnetic or charge order. Generally, repulsive Hubbard interactions suppress the charge susceptibility, but enhance the spin susceptibility  Takimoto et al. (2004); Yada and Kontani (2005); Kubo (2007); Graser et al. (2009); Liu et al. (2013); Zhang et al. (2022); Kuroki et al. (2008). Therefore, $U_{c}^{(s)}<U_{c}^{(c)}$ and we denote $U_{c}^{(s)}$ as $U_{c}$ in the bilayer subsystem in the following. When fixing $J_{H}=U/6$, we find $U_{c}\approx 1.15$ eV.
[…]
When $U<U_{c}$, the spin fluctuations can mediate SC in the bilayer subsystem, whose $T_{c}$ is related to the largest pairing eigenvalue $\lambda$ via $T_{c}\propto\omega_{D}e^{-1/\lambda}$ Graser et al. (2009), where the “Debye frequency” $\omega_{D}$ represents the typical energy scale of spin fluctuations. The pairing symmetry is determined by the corresponding eigenvector (see $\bf Methods$). Fig. 4(c) illustrates the dependence of the largest pairing eigenvalue, denoted as $\lambda$, on the interaction strength $U$ for different potential pairing symmetries. The $D_{4h}$ point group of this bilayer subsystem structure allows for several possible pairing symmetries. In Fig. 4(c), we show the three leading pairing symmetries: $s$-wave, $d$-wave, and degenerate $(p_{x},p_{y})$-wave pairings. It is clear that the $s$-wave is the leading pairing symmetry and dominates other ones. The gap function of the obtained $s$-wave pairing is shown on the FS in Fig. 4(d), which displays the $s^{\pm}$ pattern. Consequently, the $\alpha$- and $\gamma$- pockets connected by the nesting vector $\mathbf{Q}_{1}$ are distributed with the strongest pairing amplitude, with opposite gap signs. This pairing pattern is also similar to that in pressurized bulk La3Ni2O7 Luo et al. (2024); Yang et al. (2023a); Liu et al. (2023).
@@
Table 2: Hopping parameters and site energies of TB model for La5Ni3O11 under various pressures. $x$, $z$ and xzxz denotes hopping within and between $d_{x^{2}-y^{2}}$, $d_{z^{2}}$ orbitals, respectively. The abbreviations BI, BO, and SI represent intrabilayer, interbilayer, and single-layer in-plane hoppings, respectively. Only hopping integrals greater than 1 meV are listed here for clarity. All parameters are in unit of eV.
@@
The decreasing $\lambda\sim P$ relation caused by the reduced DOS shown in Fig. 5 parallels that in the bulk La3Ni2O7 Jiang et al. (2025); Huo et al. (2025); Zhang et al. (2024d), which closely matches the experimentally observed $T_{c}\sim P$ relation in that material Li et al. (2025). However, it seems at a glance that such a decreasing $\lambda\sim P$ relation conflicts with the $T_{c}\sim P$ relation in La5Ni3O11, which shows a pronounced increase up to around 20 GPa, followed by a gradual decline afterward Shi et al. (2025). The discrepancy between the $\lambda\sim P$ and the $T_{c}\sim P$ relations in La5Ni3O11 originates from that its alternating SL-BL structure brings into extremely weak IJC, which is crucial for establishing bulk superconductivity. Since superconducting pairing originates from the bilayer subsystem while global phase coherence relies on interlayer coupling, the enhancement of IJC under pressure becomes decisive, thereby resolving the seeming contradiction. In the following, we clarify this viewpoint.
Figure 6: (color online) Influence of the IJC on the pressure-dependence of the $T_{c}$. (a) Schematic of the intrinsic S–N–S Josephson Junction structure of La5Ni3O11, where the superconducting (S) layers correspond to the BL subsystem and the normal (N) layers correspond to the SL subsystem. (b) Schematic of the interlayer distance $d_{\perp}$ dependence of the hopping integral $t_{z}^{BS}$ between the BL and the SL subsystems. (c) $T_{c}/\rho_{0}$ as a function of the anisotropy parameter $\eta$ that reflects the strength of the IJC. The red dashed line marks the $\eta\approx\mathrm{1}0^{-8}$ associated with the La5Ni3O11, while the blue dashed line marks the $\eta\approx\mathrm{1}0^{-4}$ associated with the bulk La3Ni2O7. (d) Pressure-dependence of the superconducting $T_{c}$ (in unit of $\omega_{D}$) in La5Ni3O11.
As shown in Fig. 6(a), the La5Ni3O11 hosts alternating bilayer and single-layer NiO2 subsystems, where the former supports SC and the latter is a “bad metal”. This results in an intrinsic superconductor–normal metal–superconductor (S–N–S) Josephson Junction structure. In this structure, the establishment of bulk SC requires not only intra-bilayer pairing but also inter-bilayer phase coherence. Given that the intermediate single-layer subsystem is close to a Mott-insulating state, it may be approximated as an intermediate layer without itinerant electrons participating in superconductivity. Physically, the inter-bilayer phase coherence is established via combined tunneling of a Cooper pair from a superconducting bilayer to its adjacent one, i.e. the IJC, across the intervening normal-metallic single-layer. This is a second-order process with respect to the tiny inter-bilayer single-particle tunneling, and hence a fourth-order process with respect to the weak bilayer-single-layer tunneling for a single particle. As such attained IJC is extremely weak, it will experience a prominent enhancement with the increase of the pressure. In the following, we provide a more quantitative clarification for this point.
@@
In quasi-2D layered superconductors such as La5Ni3O11 and La3Ni2O7, the bulk $T_{c}$ is related to the IJC through the following relation in the limit of vanishingly weak IJC ($\eta\to 0$Kopeć and Polak (2000),
@@
Here $\rho_{0}$ indicates the intralayer phase stiffness reflecting the pairing within the bilayer subsystem, which is approximated as the pairing temperature obtained from the RPA calculation, i.e. $\omega_{D}e^{-1/\lambda}$. The anisotropy parameter $\eta$ reflects the strength of the tiny IJC, i.e. $\eta\sim J_{\perp}^{BB}$. Since $J_{\perp}^{BB}\propto(t_{z}^{BS})^{4}$, and $t_{z}^{BS}\sim 10^{-2}$, we have $\eta\sim 10^{-8}$ for La5Ni3O11, which is extremely weak. In contrast, for La3Ni2O7, since there exists no intervening normal-metallic single-layer to go across, we have $\eta\sim 10^{-4}$. Substituting this into Eq. (4) yields an estimated $T_{c}$ for La5Ni3O11. As shown in Fig. 6(c), for La5Ni3O11, the anisotropy parameter $\eta$ locates within the regime where a slight enhancement of $\eta$ will prominently enhance the $T_{c}$; while for La3Ni2O7, $\eta$ has only a minor effect on $T_{c}$. It should be noted that the $T_{c}$ here refers to the three-dimensional bulk superconducting $T_{c}$, which exhibits true zero resistance, unlike the two-dimensional Kosterlitz–Thouless transition temperature $T_{KT}$, where the resistance (or conductance) exhibits a power-law decay to zero with voltage.
Under applied pressure, the strain is proportional to the pressure, implying that $\Delta d_{\perp}\propto-P$, i.e., $d_{\perp}(P)=d_{\perp}^{0}-\kappa P$, where $d_{\perp}^{0}$ is the interlayer distance at zero pressure and $\kappa$ is the modulus constant. Therefore,
$\eta\sim e^{-4(d_{\perp}^{0}-\kappa P)/\xi}\equiv\eta_{0}e^{\alpha P}.$ (5)
@@
Since $t_{z}^{BS}$ is too small for DFT to capture its pressure dependence with sufficient accuracy, a precise calculation is challenging; therefore, we extract the parameters by fitting to experiment. For a reasonable set of parameters, e.g. $\eta_{0}\approx 10^{-8}$ and $\alpha\approx 0.3$ GPa-1, the resulting pressure dependence of $T_{c}$ shown in Fig. 6(d) indeed exhibits a dome-shaped behavior similar to that observed in experiment Shi et al. (2025).
@@ Discussion
In this work, we investigate the electronic structure and superconducting mechanism of the hybrid RP nickelate La5Ni3O11, composed of alternating single-layer and bilayer subsystems. Based on DFT and RPA calculations, we find that SC predominantly emerges in the bilayer subsystem with an $s^{\pm}$-wave pairing symmetry, closely resembling that in pressurized bulk La3Ni2O7. We further show that the pressure dependence of the superconducting $T_{c}$ in La5Ni3O11 is governed not only by the intralayer pairing but also by the IJC, which leads to a dome-like $T_{c}(P)$ behavior consistent with experimental observations. Although our analysis reproduces this dome-like trend, this model should be regarded as one possible channel to account for the observed $T_{c}$ of La5Ni3O11, while other channels may also be relevant.
Figure 7: (color online) The distributions of the leading $d_{xy}$-wave pairing gap function with $\lambda=0.435$ at $U=0.5$ eV, $J_{H}=1/6U$.
Recently, a full monolayer–bilayer RPA study was reported in Ref. Zhang et al. (2025b). We therefore also performed a full monolayer–bilayer calculation, as shown in Fig. 7.
@@ Methods
DFT calculations were performed by Vienna ab initio simulation package (VASP) (Kresse and Furthmuller, 1996a, b), in which the projector augmented wave (PAW) (Blochl, 1994; Kresse and Joubert, 1999) method with a 600 eV plane-wave cutoff is applied. The generalized gradient approximation (GGA) of Perdew-Burke-Ernzerhof (PBE) form exchange correlation potential is adopted  (Perdew et al., 1996). The convergence criterion of force was set to be 0.001 eV/Å and total energy convergence criterion was set to be $10^{-7}$ eV. A $\Gamma$-centered $19\times 19\times 5$ Monkhorst Pack k-mesh grid is used for primitive cell of $P4/mmm$ phase and a $\Gamma$-centered $14\times 14\times 5$ Monkhorst Pack k-mesh grid for primitive cell of $Cmmm$ phase. In structural relaxations, the lattice constants of $P4/mmmm$ phase we use are $a=3.77$ Å and $c=16.20$ Å. DFT calculations reveal that the corresponding pressure is 12 GPa, where La5Ni3O11 enters superconducting state  Shi et al. (2025). An effective Hubbard U parameter $U=3.5$ eV was employed to account for the correlation effects of 3d electrons in Ni atoms (Yang et al., 2024c; Dudarev et al., 1998). To obtain the projected TB models, we further performed Wannier downfolding as implemented in WANNIER90 (Mostofi et al., 2008) package, in which the good convergences were reached.
@@
$\left.\times c_{s}^{\dagger}(\bm{k}_{2}+\bm{k},0)c_{t}(\bm{k}_{2},0)\right\rangle_{0},$ (7)
@@
[…]
@@
Note that there is a critical interaction strength $U_{c}^{(s,c)}$ for spin and charge respectively. When $U>U_{c}^{(s,c)}$, the denominator matrix in Eq. Pairing mechanism and superconductivity in pressurized La5Ni3O11 will have zero eigenvalues for some $\mathrm{q}$ and the renormalized spin or charge susceptibility diverges there, which invalidates the RPA treatment Takimoto et al. (2004); Yada and Kontani (2005); Kubo (2007); Kuroki et al. (2008); Graser et al. (2009); Liu et al. (2013); Wu et al. (2014); Ma et al. (2014); Zhang et al. (2024b). This divergence of spin susceptibility for $U>U_{c}^{(s)}$ implies magnetic order, while that of the charge susceptibility for $U>U_{c}^{(c)}$ implies charge order.
@@
Here $\alpha$$\beta$ is band indice. And effective pairing interaction vertex $V^{\alpha\beta}(\mathbf{k},\mathbf{q})$ has the form:
@@
Here $\xi^{\alpha}\left(\mathbf{k}\right)$ are the eigenvector corresponding to $\alpha$-th eigenvalue (relative to the chemical potential $\mu_{c}$) of the TB Hamiltonian matrices. Within the mean-field approximation, one can derive a self-consistent equation for the pairing gap, which when linearized near $T_{c}$ Takimoto et al. (2004); Yada and Kontani (2005); Kubo (2007); Kuroki et al. (2008); Graser et al. (2009); Liu et al. (2013); Wu et al. (2014); Ma et al. (2014); Zhang et al. (2024b) yields the linearized gap equation (12):
@@
Here $\beta_{c}$$\frac{1}{\mathrm{k_{B}}T_{c}}$ represents the critical temperature of SC. Choosing a thin energy shell near the Fermi level, Eq. (12) becomes the eigenvalue problem of the effective interaction matrix $V^{\alpha\beta}$, which determines the $T_{c}$ and the leading pairing symmetry of the system. To be specific, $\Delta_{\alpha}(\mathbf{k})$ represents the relative gap function on the $\alpha$-th FS patches near $T_{c}$, and eigenvalue $\lambda$ is related to $T_{c}$ through $\lambda^{-1}=\mathrm{ln}(1.13\frac{\hbar\omega_{D}}{k_{\mathrm{B}}}T_{c})$. The leading pairing symmetry is determined by the largest eigenvalue $\lambda$ of Eq. (13):
$\frac{1}{(2\pi)^{2}}\sum_{\beta}\oint_{FS}dq_{\|}\frac{V^{\alpha\beta}(\mathbf{k},\mathbf{q})}{v^{\beta}_{F}(\mathbf{q})}\Delta_{\beta}(\mathbf{q})=\lambda\Delta_{\alpha}(\mathbf{k}).$ (13)